OpenStudy (anonymous):

x/(x^2 - 2) = 2/(3x + 4) solve the proportion

OpenStudy (amistre64):

we can "cross multiply" and then collect the terms to one side and solve a quadratic i believe

OpenStudy (anonymous):

uh ok can you do that for me I'm taking a pretest so i haven't learned this yet

OpenStudy (amistre64):

the pretest is the place TO learn it :)

OpenStudy (anonymous):

please i need a good grade on this

OpenStudy (amistre64):

\[\frac{a}{b}=\frac{c}{d}\to \ ad=bc\]

OpenStudy (amistre64):

thats the essense of a cross multiply

OpenStudy (anonymous):

um ok i'll try that so x(3x+4)= 2(x^2 - 2)

OpenStudy (amistre64):

x(3x+4) = 2(x^2 - 2) i agree so far; now lets distribute thru on each side ...

OpenStudy (amistre64):

\[x(a+b)\to \ ax+bx\] is distribution

OpenStudy (anonymous):

ok so 4x + 4x = 2x^2 - 4 ???

OpenStudy (amistre64):

almost; that 3xx kinda snuck away from you :) 3xx+4x = 2x^2 - 4 3x^2 +4x = 2x^2 - 4 now just move all the terms to one side using the appropriate adds and subtracts

OpenStudy (amistre64):

3x^2 +4x = 2x^2 - 4 -2x^2 +4 -2x^2+4 ------------------------ x^2 +4x +4 = 0

OpenStudy (anonymous):

now i do something with like quadratic equation right?

OpenStudy (amistre64):

yes, and depending on how familiar you are with them this can be easy or tricky ...

OpenStudy (amistre64):

it factors nicely tho if your comfy with that

OpenStudy (anonymous):

yeah i don't really remember how to do it

OpenStudy (amistre64):

yeah, it takes practice :) that end is a 4; we need the factors of 4: 1*4 2*2 the middle number is a 4; we need a set of factors from that first part that ADD up to 4 1+4 = 5 ; not it 2+2 = 4 ; that one will do just fine split the quadratic into its component parts and fill in our numbers (x+2) (x+2) = 0 is then what we solve

OpenStudy (amistre64):

when x=-2 we get: (-2+2) (-2+2) = 0 0(0) = 0 0=0 so when x=-2 is the solution to all our problems .... in this case :)

OpenStudy (anonymous):

oh yesss now i remember so x= 2 and x = -2 ???

OpenStudy (anonymous):

or is it just x= -2

OpenStudy (amistre64):

just x=-2 ... thats the only thing that will fit

OpenStudy (amistre64):

if we try x=2 (2+2)(2+2)=0 4(4) = 0 16=0 just doesnt fit

OpenStudy (anonymous):

okay thanks can u answer my other question. its the last question i asked.

OpenStudy (amistre64):

i can give it a shot